Update 26 October 2016
Field on LiH programme started
Data taking for the field-on LiH scattering programme started on Monday 24Oct16.
The attached plots show the difference between the momentum reconstructed in the upstream and downstream trackers for 140MeV/c and 170MeV/c beams. The distributions show that the momentum is decreased by roughly 15MeV/c as the beam passes through the absorber. The size of this momentum loss is in line with expectations.
Both spectrometers were ramped to the currents required by the “6.1.4” scattering configuration late on Sunday 25Oct16.
The FC was ramped to current over night Sunday/Monday. During the day on Monday 24Oct16, ISIS recovered from the issues they had experienced in the vacuum and RF systems. MICE began taking data as soon as ISIS was ready, data taking continues.
By tomorrow morning, we anticipate that substantial scattering data sets will have been accumulated at 140MeV/c, 170MeV/c, 200MeV/c and 240MeV/c. Therefore, in preparation for the November/December Cycle, data will be taken tomorrow with various diffuser settings. This data will allow us to verify our understanding of the effect of the diffuser on the beam; an exercise that we have not performed before.
This is a big step for our physics programme and is the result of a lot of hard work by many, particularly over the past two weeks during which the absorber was installed and gremlins in the solenoid power systems were flushed out.
To recognise what has been achieved, and to initiate the discussion of the programme to be carried out in the next User Cycle, you are cordially invited to a MICE end-of-cycle briefing which will be held in R9 at 12:00 on Friday 28Oct16.
With very best wishes …